Bobby traveled a distance of 50x + 150 miles at 50 miles per hour on the Gulf Freeway. Find a binomial that represents the time that he traveled.
Answer by venugopalramana(3286)

Bobby traveled a distance of 50x + 150 miles at 50 miles per hour on the Gulf Freeway. Find a binomial that represents the time that he traveled.
time*speed=distance...or...time=distance/speed
=(50x+150)/50=(50x/50)+(150/50)=x+3
time of travel =x+3...which is a binomial